For each of the following, compute the future value (Do not round intermediate calculations and round your final answers to 2 decimal places. (e.g., 32.16)):

       
    

Present Value Years Interest Rate Future Value
$ 2,450 10      14 % $   
9,152 8      8   
80,355 15      13   
187,796 6      5   
Answer #1

media%2F6da%2F6dacc844-c918-4d21-9471-2f

PV is the present value

t is the number of compounding period

i is the interest rate for that period


a.)FV=2450(1.14)^10=$9082.69

b.)FV=9152(1.08)^8=$16939.71

c.)FV=80355(1.13)^15=$502561.90

d.)FV=187796(1.05)^6=$251664.60